# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ 5.2 Qualitative Findings on Non-Algorithmic Targeting
“Student wants to transfer to GT [Georgia Tech]. Says they has conditional acceptance after 30 credit hours. I explained that not all CS courses transfer to GT and they need to find out what transfers. For spring, they want to take 18 credits to ensure they will reach 30 credit hours at the end of spring.”
While it is difficult to confirm these factors are algorithmically unpredictable from basic administrative data or directly influence outcomes, the complexity of the reasoning behind deciding on an intervention and the variety of the reasons provided for those actions make this plausible in many cases.
The process of choosing how to act on non-algorithmic information also involves expertise. One piece of information could support multiple paths of action. The comments show how advisors exercise discretion to select an appropriate intervention from multiple options, based on the information they gather, attending to subtle details to decide the appropriate course of action. For instance, when a student is struggling with their coursework, advisors sometimes do not just recommend that students correct their schedule, as in the above cases, but instead warn the student about more dire progression issues (Intervention 4), based on their judgment of the severity of the problem. One advisor, for example, suggests a student reconsider their enrollment in the pre-nursing program based on their reported difficulty with their science courses:
“Student in pre-nursing major stated that math was giving them trouble. Without being discouraging, I explained the GPA requirements.”
Moreover, within interventions, the distribution of kinds of information that can persuade the advisor to act often has a long tail: in most cases, there are many different kinds of information that can lead an advisor to take a particular action, making it difficult to improve a targeting algorithm by simple augmentations of the dataset and data collection process. Table 3 demonstrates the diversity of the reasons advisors cite for taking all the significant interventions we analyzed. When advising a schedule correction (Intervention 2), for instance, advisors consider matters including student’s course preferences, performance, financial aid requirements, projected course loads, prior placement exam performance, and credits earned elsewhere (which often have not been recorded in administrative data at the time of their conversation).
Finally, our analysis reveals that these patterns of plausibly non-algorithmic, diverse information supporting discretionary action are largely unique to the significant interventions (Appendix A.5.4).
(B) Evidence for targeted action supporting at-risk students.
Among the significant interventions that predict negative future outcomes, advisors take actions based on non-algorithmic information indicating potential risk factors to the student’s success.
Consider the intervention of recommending a student schedule an appointment (Intervention 16). Most often, advisors take this action for students who fail to schedule appointments or miss their scheduled appointments, an intuitive risk factor for poor outcomes:
“I called student because no appointment yet.”
@@
“They are currently living with friends and family. Apparently their mom is trying to figure out a new living situation and because of such, they do not have a permanent place to call home.”
@@
Table 3 shows additional plausible risk factors, like failure to enroll or poor course performance, guiding advisors’ requests to schedule an appointment and the other interventions that predict poor outcomes, recommending a schedule correction (Intervention 2), and warning about progression issues (Intervention 4). This supports the hypothesis that these actions negatively predict outcomes because advisors are performing these actions to support at-risk students that go unidentified by risk prediction algorithms based on limited administrative data, rather than that advisors are performing actions that have a negative effect on the students.
(C) Some interventions are confounded by nuisance factors.
While there is a pattern of interventions showing signs of non-algorithmic targeting, some also show signs of a confounding effect, in which advisors perform interventions after students explicitly request they do. One such case is recommending a schedule (Intervention 17), which positively predicts next-term persistence ($p$=0.039), as students often seek scheduling advice for the next semester (Appendix A.5.3):
@@
“Student wanted to talk about college career.”
@@
Such student requests confound the relationship between actions and outcomes with a factor that is arguably not non-algorithmic information. A student requesting schedule recommendations for next semester from their advisor likely also intends to enroll regardless of the advisor’s particular recommendations. Similar confounding also occurs in comments related to schedule corrections (Intervention 2), progression issues (Intervention 4), being advised to schedule an appointment (Intervention 16), and developing the degree planner (Intervention 20) as the counts in row 12 of Table 3 (student-initiated interventions) demonstrate.
Significance on ExpertTest*: $\times$ $\times$ $\times$
Reason Category Int 2 Int 4 Int 5 Int 9 Int 10 Int 16 Int 20 Category Sum
1 Course Credits and Placement 3 4 5 - - - 1 12
2 Course Problem 11 6 6 - - 6 2 29
3 Degree Progress - 8 3 - - - 2 11
4 Documents - - - 5 - - - 5
5 Fin. Registration Problem - - - 52 - 4 - 56
6 Load Management 18 - 1 - - - - 19
7 Major Choice - 8 1 - - 18 - 27
8 Non-fin. Registration Problem - - - - 3 - - 3
9 Personal Circumstances - 2 - 1 - 5 1 8
10 Low Engagement - - - - - 109 - 109
11 Scholarship Process 2 - - 6 - 1 - 9
12 Student-Initiated 3 - - 1 - 8 1 12
13 Transfer/Dropout - 3 - - - 3 - 6
14 Unregistered - - - - - 12 - 12
Table 3: Counts of specified reason categories for each intervention type mentioned during advising meetings, indicating how frequently each reason was associated with a particular intervention, for Intervention 2 (Correct Schedule), Intervention 4 (Progression Issues), Intervention 5 (Adjust Course Balance), Intervention 9 (Address Financial Issues), Intervention 10 (Address Non-Financial Collegiate Requirements), Intervention 16 (Schedule Appointment), and Intervention 20 (Develop Degree Planner). Results for Intervention 17 (Recommend Schedule) are presented separately in Appendix A.5.3 since we look for signs of non-algorithmic information in schedule customizations rather than reasons to recommend a schedule.
@@ 6 Discussion
Our study of the MAAPS intervention shines a light on the central role of discretionary action by advisors in the use of an algorithmic decision support for advising. This role for discretion raises key questions for future algorithmic systems, introduces new challenges for existing methodology, and suggests the importance of reflecting on the normative role of discretion in institutional decision making.
6.1 Future Algorithmic Decision Systems: Design, Implementation, and Evaluation
The central role of discretionary decision making in the MAAPS study has important implications for the design, implementation, and evaluation of algorithmic decision aids, in advising and beyond.
6.1.1 Designing for Human Discretion
Many examples of non-algorithmic information used by MAAPS advisors point to sources of data that are inherently difficult, if not impossible, to encode quantitatively. Advisors, for instance, frequently reference psychological constructs such as student engagement and motivation to justify their interventions, which are notoriously hard to measure [22, 23], but can be observed by advisors through conversation. Moreover, consistent with findings from other domains where human discretion helps compensate for flawed or incomplete data [15], advisors in the MAAPS experiment also draw on non-algorithmic information to fill gaps in administrative data that risk prediction tools use—such as missing course credit information or delays in placement test results.
Even if technologies like language models can gather some of this information from students, interpersonal relationships may still be essential for eliciting sensitive information and for providing meaningful support [56]. Moreover, expanding structured data collection about students also raises privacy concerns that discretionary discussion with advisors can avoid.
Just as there are obstacles to encoding this information in an algorithmic system, these systems may also struggle to learn how to use this kind of information appropriately. Advisors often construct elaborate causal narratives about a student’s circumstances and likely behavior to guide their interventions. Even routine suggestions—such as advising a student to drop a course—can require intricate causal reasoning: how past coursework may affect future scheduling, why a student wants to withdraw, and, if the student plans to retake the course, whether their obstacles to succeeding will apply in the future. [^26] For example: “The student wanted to withdraw from their economics class, as their first exam was a 40%. They did admit they just didn’t study, and they understand the importance now. Because they brought in AP credit, they can withdraw and still be on track. But I made it clear that they cannot withdraw moving forward or else they’ll be forced to take summer classes to stay on track. They understand.” While human advisors naturally draw on this kind of causal reasoning—as has been observed across decision-making contexts [37]—current algorithmic systems are not built to learn these causal relationships [56]. [^27] Such as the Navigate tool [20] or the student-support chatbots used at GSU [48]. Moreover, our qualitative analysis of non-algorithmic information also underscores the importance of discretion in handling rare cases, a task at which even state-of-the-art language models often struggle, both in general and in advising contexts specifically [36, 41]. Rather than aiming to replicate human discretion, future systems might be more effective if designed to support and enhance human expertise [45]. This could be done by training new advisors to develop expertise by embedding expert strategies in large language models [70], or by automating routine administrative actions that do not require discretionary expertise (discovered through techniques such as those in the current work), thereby allowing advisors to focus on building relationships with students and exercising judgment where needed.
6.1.2 Implementation Challenges of Generalizing and Scaling
The MAAPS study’s positive treatment effects for graduation and credit accumulation at GSU shows the promise of algorithmic decision aids to improve educational outcomes. However, as with many educational interventions, MAAPS faces two key challenges to external validity: generalizing the intervention to new populations and scaling it across institutional contexts [39, 66, 34, 43]. These challenges are exacerbated in studies of algorithmic decision aids where human decision making plays a central role, like the MAAPS study. In such cases, positive outcomes cannot be neatly attributed to the decision aid alone; they also depend on advisors’ ability to skillfully mediate between the system and the students. Thus, scaling this intervention requires not only scaling the algorithmic component but also hiring similarly skilled human advisors who can use the tool. These discretionary interactions are difficult to replicate in new populations or standardize across institutional contexts, complicating both generalization and scalability.
Generalizing the intervention is complicated by both institutional culture and student demographics. At GSU, the MAAPS protocol was a natural extension of the schools’ existing advising approach, which likely facilitated its adoption by advisors. By contrast, advisors at other institutions may have been less receptive to the protocol, particularly if it departs from their advising norms [60]. Moreover, at GSU, a plurality of students are African American [17], which may shape how students experience predictive interventions. The risk of racialized stigma to being flagged as high-risk by an algorithm [69] may increase significantly in contexts where structurally disadvantaged groups are numerical minorities, especially if demographic characteristics are included as inputs in the predictive model (as they are not at GSU) [20].
Similarly, implementation challenges can blunt the effectiveness of the MAAPS intervention when scaled. Institutions that do not fully adopt the MAAPS advising protocol are unlikely to realize the same benefits from algorithmic decision aids as those observed at GSU. The MAAPS study itself illustrates this point: other participating institutions in the study struggled to implement core components of the intervention, such as assigning students to primary advisors from a centralized advising office rather than across academic departments [1]. None of these institutions achieved the same positive treatment effects as GSU, which the study authors partly attribute to these implementation challenges [60].
Another likely contributor to the success of the MAAPS intervention, and to predictive analytics more broadly at GSU, is the university’s investment in increasing the capacity of the human decision-makers. Prior work has shown that increased capacity in human outreach programs leads to larger effect sizes and higher power in RCTs [11], supported by the results of the RCT at GSU versus the other participating institutions. As part of its advising strategy, GSU reduced its student-to-advisor ratio from 700:1 to 300:1 [58]. Within the MAAPS intervention treatment arm, this ratio dropped to approximately 150:1 [60]. Many of the other institutions participating in the MAAPS were unable to consistently maintain this ratio due to advisor retention problems [60]. These numbers suggest that the effectiveness of decision aids may depend on the institutional ability to support and retain human advisors. Schools that adopt algorithmic systems as cost-saving measures—placing human decision-makers in resource-constrained situations or pressuring them to follow algorithmic recommendations [37]—are unlikely to replicate positive outcomes.
6.1.3 Evaluating Algorithms: Human Decision-makers as a Study Condition
The importance of human discretion to advising also has implications for designing future studies of decision support systems. Without understanding how the interaction between humans and algorithmic tools drives positive effects, replicating or scaling those effects of an implementation is difficult. Indeed, randomized field trials may depend on factors about the human’s cognitive response that could change outside the experimental setting—for instance, an increase in the false positive rate of algorithmic decision aid might influence how much human decisionmakers rely on it [55]. Thus, as part of developing randomized trials of real-world algorithm deployments [33], we argue it is important to develop techniques that build a thorough understanding of how the human decision-maker reacts to the decision aid.
Some researchers have attempted to study human discretion through controlled experiments with crowdworkers [26, 14, 54, 27], but such studies often lack external validity, failing to capture the knowledge of real decisionmakers and the complexity of institutional contexts [47]. By contrast, randomized field experiments can look at the kinds of discretion that arise in the wild, using a mix of quantitative and qualitative methods. Randomly assigning advisors to students in treatment, for instance, could help disentangle the effects of human discretion from those of the algorithmic tool itself. Additionally, studies of algorithmic decision aids should collect data—both structured and unstructured—on how humans interpret and act on algorithmic recommendations. The MAAPS study offers a promising model in its documentation of advisor actions and rich comments on meetings; this detailed picture of advisor decision making allows us to take a uniquely granular look at factors influencing advisor decision making in this paper. This data could be further improved by collecting records of the algorithmic outputs themselves and descriptions of how advisors used them.
6.2 Methodological and Practical Limitations in Capturing Human Expertise
Our study of discretionary decision making within the MAAPS dataset has two key limitations, both of which highlight broader challenges for future research on discretionary algorithm-supported decision making. First, our theoretical framework is designed for single-stage decision making, but collegiate advising (as in the MAAPS experiment) is a multistage process. Advisors work with students across multiple advising meetings, helping them achieve long-term academic and personal goals. As a result, the definition of expertise presented here (and the statistical condition used for the ExpertTest*) does not capture forms of human expertise that may not be immediately observable at one stage of decision making. For instance, an expanded notion of expertise might capture trajectories of actions leading to good outcomes. This raises a broader challenge in understanding discretionary decision making when the effects of decisions may be cumulative or take a long time to materialize.
Second, our study is limited by the lack of access to algorithmic outputs, such as the risk scores used in the MAAPS intervention. Direct access would enable us to understand advisor expertise relative to actual risk assessments. While our use of the ExpertTest* allows us to assess whether advisors target interventions for any hypothetical algorithm, it constrains the types of analyses we can perform. For instance, we did not incorporate recurring time-series data (Appendix A.1.2), as the test is not designed for such data. Future research could build on this work by collecting algorithmic outputs and exploring alternative dependence tests.
Finally, nuisance confounding, such as explicit student requests influencing advisor actions, hinders the ability of our statistical test to indicate potential expertly targeted action. This limitation underscores the inherent hazards of using observational data for studying causal questions and the importance of creating evaluation techniques that can build a causal understanding of discretion in decision making through controlled experimentation.
@@ 6.2 Methodological and Practical Limitations in Capturing Human Expertise
6.3 Functional and Normative Roles of Human Discretion in Institutional Decision Making

### v3

@@ 11 Unregistered - - 12 - -
“Student wants to transfer to GT [Georgia Tech]. Says they have conditional acceptance after 30 credit hours. I explained that not all CS courses transfer to GT and they need to find out what transfers. For spring, they want to take 18 credits to ensure they will the reach 30 credit hours.”
All of these comments demonstrate examples where non-algorithmic context informs advisor actions. For example, in the third comment, the context $\mathbf{U_{t}}$ is the student’s requirements to transfer to Georgia Tech, and the $A_{t}$ is designing a schedule to contain 18 hours.
In many cases, advisors incorporate context $\mathbf{U_{t}}$ that function as non-algorithmic risk factors, signaling that a student may be at risk of worse outcomes. Consider recommending the student to schedule a meeting (Intervention 16), which is negatively associated with a high GPA ($p=0.039$). Most often, advisors take this action for students because of the context $\mathbf{U_{t}}$ that these students failed to schedule appointments or miss their scheduled appointments, an intuitive risk factor for poor outcomes:
@@
“They are currently living with friends and family. [Redacted], and because of such, they do not have a permanent place to call home.”
@@
The targeting based on non-algorithmic risk factors $\mathbf{U_{t}}$ may explain why Intervention 16 is negatively associated with GPA. The other two significant interventions associated with a low GPA—recommending a schedule correction (Intervention 2) ($p=0.049$) and advising about progression issues (Intervention 4) (p0.049)(p-0.049)—show a similar pattern of targeting based on risk factors such as course performance, degree progress, and challenging personal circumstances, as Table 3, which includes systematic counts of reasons advisors take each intervention, demonstrates.
Finally, our qualitative analysis suggests it would be difficult to capture all this rich context without human expertise. In particular, dataset augmentation and algorithmic enhancements will likely be insufficient, for two reasons. First, for any given intervention, the distribution of kinds of information that can persuade the advisor to act has a long tail: in most cases, there is a broad spectrum of information that can lead an advisor to take a particular action. Table 3 demonstrates the diversity of the reasons advisors cite for taking significant interventions. When advising a schedule correction (Intervention 2), for instance, advisors consider matters including student’s course preferences, performance, financial aid requirements, projected course or extracurricular loads, prior placement exam performance, and credits earned elsewhere (which often have not been recorded in administrative data at the time of their conversation).
Second, subtle nuances in context, which may be flattened or ignored by tabularization, are consequential to the actions advisors take. Consider the following pair of examples for schedule corrections (Intervention 2):
“Student needs to add more credits to fall. Offered [MATH XXX]. Student didn’t want to take [MATH XXX] because they said they needed a ‘break’ from math. However, this will help the student stay on track, so I strongly encouraged them to take the class in the fall.”
These two situations, in which a student wants to postpone a math course, appear similar at first but differ in important ways that lead to different advisor recommendations. The first student has a clear plan to make up the course and a track record of success in summer courses. The second student does not offer a clear reason for postponing the math course and needs more courses, suggesting they would not benefit in the same way from a postponement recommendation. Adequate dataset augmentation may require taking such complex signals of student motivation into account.
6.2.2 Alternative Explanations: Evidence of Confounding
Conditional dependence is only a necessary condition for expertise. As summarized in §5.4, there are three alternative explanations for significant results.We ask whether our qualitative data support any of these alternative explanations, and find evidence of nuisance confounding, especially in recommend schedule (Intervention 17).
Some Interventions are Confounded by Nuisance Factors.
While significant interventions show signs of non-algorithmic targeting, some additionally show signs of a confounding effect, in which advisors perform interventions after students explicitly request they do. [^26] We note the presence of confounding does not rule out the simultaneous presence of expertise.
The clearest such case is recommending a schedule (Intervention 17), which positively predicts next-term persistence (pp=0.039). Consider the following examples:
@@ 6.2.2 Alternative Explanations: Evidence of Confounding
Such student requests, of which there are 243 examples, confound the relationship between actions and outcomes: A student requesting schedule recommendations for next semester from their advisor likely also intends to enroll regardless of the advisor’s particular recommendations. At a much smaller scale, similar confounding also occurs in comments related to schedule corrections (Intervention 2), progression issues (Intervention 4), and being advised to schedule an appointment (Intervention 16), as the counts in row 9 of Table 3 (student-initiated interventions) demonstrate.
Evidence Against Absent Targeting for Most Interventions.
For meeting comments associated with four of the five significant interventions, we see comments where advisors incorporate non-algorithmic information into their recommendations. This provides evidence against the absence of targeting altogether. For the remaining significant intervention, advise through in-person development of the degree map or planner (Intervention 20), we do not have enough data to determine whether the action is non-algorithmically targeted. Advisors only explicitly describe their reason for developing the degree planner in 6 out of 284 comments associated with meetings in which the intervention was performed (Appendix Table 16).
Inconclusive Evidence about Ineffective Action or Lack of Heterogeneous Effects.
The qualitative analysis does not provide adequate evidence about ineffective action or lack of heterogeneous effects as an alternative explanation. Comments cannot reveal how advisor interventions and non-algorithmic information causally affect outcomes; they can only tell us how advisors choose actions to take. We note that many of the above instances of targeting that we identified plausibly are helpful to students and use non-algorithmic context relevant to the effect of the action. For example, take the case of the student with a health condition whose schedule must be adjusted so their classes are in close proximity. It is plausible that this schedule customization helped the student, and that the incorporation of context about their medical condition was important to the intervention’s success.
6.3 Qualitative Findings on Advisor Styles
Our exploratory qualitative analysis of advisors’ comments reveals notable differences in their advising styles. Out of three advisors with a substantial numbers of comments available for analysis, while Advisor 1 and 2’s comments show what we term an academic style (Appendix B.2) focusing on topics like schedule planning and course performance, Advisor 3’s comment reflect what we call a holistic style of broader engagement with the student, addressing not only academic planning, but also discussion of non-academic factors like career goals and family circumstances, observations about student personality, and expressions of empathy for student circumstances.
Figure 2: Fraction of meeting comments coded with academic, non-academic, and relationship codes (Appendix B.1), broken down by advisor. Advisor 3’s comments more often bring up non-academic topics and show more signs of building an advising relationship with students. Note that fractions for different code categories within each advisor may not sum to 1 as multiple codes may be applied to one meeting.
This difference is visualized in the distribution of codes applied to their meetings (Figure 2), which shows that Advisor 3’s comments display a more even balance of academic, non-academic (discussions of topics like career aspirations, finances, or family circumstances), and relationship-building (expressions of empathy or observations about students’ personalities) codes. A full description of the codes is included in Appendix B.1.
Association with Graduation Outcomes.
To understand whether this difference in advising style is consequential, we use MAAPS trial data to look for heterogeneous treatment effects of the MAAPS intervention across advisors on four-year graduation (Appendix C.1). While we do not find statistically significant differences in effect size, logistic and linear regression models produce a consistent ranking of advisor-specific coefficients with Advisor 3 at the top (Appendix C.2). This ordering provides preliminary quantitative evidence that the holistic advising style may be more effective than the academic one.
We hypothesize that this heterogeneity may be partly explained by the more holistic advising style enabling Advisor 3 to expertly target student support better than the academic style. For example, one student confides in Advisor 3 about their mental health:
“Student is very anxious about GSU, college and career in general … They are worried they won’t be successful and that they won’t find a job and that they’ll want to change their major. I asked them to look up [people in their profession] and see what they did to get their jobs. The student is worried they won’t get exposure to their field, so I challenged them to get exposure. They did ask about the counseling center.”
The emotional connection allows Advisor 3 to learn important non-algorithmic information $\mathbf{U}_{t}$ (student anxiety) to take action $A_{t}$ (in this case, helping the student research their chosen career path and recommending mental health resources) that is plausibly important to improving the student’s outcomes.
In another example, Advisor 3 plumbs a student’s deeper motivations for pursuing a nursing degree:
“Student is in the pre-nursing program, but they stated that their math and chemistry prerequisites were giving them trouble. I explained the GPA requirements. They want to help people, but I challenged them to think of why and how.”
The comment shows Advisor 3 incorporating knowledge of the student’s underlying altruistic motivations into their recommendation to explore other, more academically feasible paths that fulfill this aim.
Advisor # Comments Average JSI
Advisor 1 770 0.5960
Advisor 2 364 0.7694
Advisor 3 360 0.4688
Table 4: Average pairwise Jaccard Similarity Index (JSI) on qualitative codes applied in the meetings for each advisor.
Finally, echoing the finding that there is a “long tail” of diverse non-algorithmic context that advisors use to support students, Advisor 3 shows more diversity in the topics covered across different meetings with students, suggesting that their meetings are more tailored to students’ unique circumstances. We measure across-meeting diversity by taking the average Jaccard Similarity Index (JSI) between codes applied to all possible pairs of each advisor’s meeting comments, with a higher average JSI indicating more similarity between that advisor’s meeting comments. [^27] For a pair of meeting comments, the JSI is defined as the ratio of the intersection to the union of the codes applied to each meeting. The maximum JSI of 1 indicates entirely overlapping codes, while the minimum JSI of 0 indicates entirely disjoint codes. We exclude comments with no codes. Table 4 shows Advisor 3 has a notably lower average JSI.
Together, these findings suggest the potential importance of advisor style in effectively targeting student support, illuminating an interesting direction for future research. It is also important to note the potential for unmeasured effects of advising style, such as student satisfaction or self-efficacy.
@@ 7 Discussion
Our study of the MAAPS intervention shines a light on the central role of advisor discretion in algorithm-supported advising. Combining quantitative and qualitative analysis, we provide diverse, suggestive evidence that discretionary action by advisors helps students achieve better educational outcomes. In this section, we discuss four implications of this finding for the design, evaluation, implementation, and normative aspects of algorithmic decision support, in college advising and beyond:
  • Algorithmic decision aids should support, not replace, advisors’ discretionary judgment (§7.1).
  • Evaluations of algorithmic decision aids that provide insight into how human discretion affects decision making “in the wild” are needed (§7.2).
  • Where human discretion is important to an algorithmic decision aid’s success, impacts of deployment may fail to generalize or scale (§7.3).
@@
  • It is important to consider both the practical and normative roles of human discretion before automating decision making (§7.4).
7.1 Designing Algorithmic Aids for Human Expertise
Much of the research on the design of algorithmic decision aids ignores the role of human decision makers, and a recent literature review finds that algorithm design in higher education is no exception to this trend [60]. Our work suggests this is a missed opportunity. When the exercise of human discretion can improve student outcomes, algorithmic aids can be made more effective by being explicitly designed to support and enhance, rather than replace, human expertise [53, 40].
There are many ways to incorporate human discretion in the design of algorithmic tools. In addition to using human-centered design principles [60], algorithm designers could draw on the literature on complementarity to optimize for joint human-algorithmic performance [26]. New advisors could be trained by large language models with embedded expert strategies [90]. Algorithmic tools could be used to automate routine administrative actions that do not require expertise, freeing advisors’ time to focus on more difficult tasks [85]. And algorithms can be used to suggest “tips” that address bottlenecks in advisor decision making [14].
Our study also identifies significant disadvantages to the increasingly frequent proposal of replacing human advisors with automation. In particular, it is difficult to capture all of the non-algorithmic context used in college advising through increased data collection or feature extraction from unstructured data (as in e.g., [89]). Much of the non-algorithmic information used by MAAPS advisors, such as the psychological constructs of student engagement and motivation, is difficult to encode quantitatively [32, 33]. Advisors also use non-algorithmic context to address data gaps, such as missing course credit information or delays in placement test results, echoing a pattern in other domains [20]. While chatbots could gather up-to-date information in principle, interpersonal relationships and trust are essential to obtaining information about sensitive topics and for providing meaningful support [73]. Finally, expanded structured data collection about students can act as a form of surveillance [21], with risks like “anticipatory conformity,” students altering their behavior to avoid being algorithmically flagged [59].
7.2 Evaluating Discretion in the Wild
The importance of human discretion to the MAAPS experiment also has implications for evaluating decision support systems. Our results highlight the need for developing experimental methodology that allows us to understand the role of discretion in the success of algorithmic aids.
Some researchers have attempted to study human discretion through controlled experiments with crowdworkers [36, 19, 71, 37], but such studies often lack external validity, failing to capture the knowledge of real decision makers and the complexity of institutional contexts [57, 42]. Thus, in experimental evaluations of real-world algorithm deployments [43], it is important to develop techniques that build a thorough understanding of how the human decision maker interacts with the decision aid, and how that interaction affects outcomes [72]. For example, experimenters can treat the human decision maker, non-algorithmic external context, and the algorithmic aid as study conditions, experimentally manipulating these conditions in order to disentangle how they interact. When manipulation is impractical, field experiments can still collect structured or qualitative data to build an understanding of this interaction.
Many of the limitations of the present work offer additional future directions for developing the methodology of studying discretion in the wild. One limitation of our framework is its focus on single-stage decision making, while collegiate advising (as in the MAAPS experiment) is a multistage process. An interesting future direction might be to generalize the framework to long-term outcomes and timeseries data about students, and design studies that can identify long-range effects. Another limitation of our study is that the MAAPS dataset lacks risk score records and descriptions of how advisors used the algorithm. Future studies could collect these data to enable a situated understanding of how algorithmic tools are used.
7.3 Implementation Challenges: Generalization and Scaling
Even when algorithm-assisted interventions are designed and evaluated with human decision makers in mind, they face two challenges common to many educational interventions: generalizing the intervention to new populations and scaling it across institutional contexts [49, 84, 45, 52]. When discretion plays a role in intervention outcomes, the success of an intervention also depends on the success of the human decision maker, which may not generalize or scale.
For example, when generalizing interventions to new institutional settings, human decision makers may face structural barriers to their success. The MAAPS study exemplifies this pattern: Unlike GSU, other participating institutions in the MAAPS study did not implement core components of the intervention, such as assigning students to primary advisors from a centralized advising office rather than across academic departments [3]. No other institutions achieved the same positive treatment effects as GSU, which the study authors partly attribute to this implementation challenge [77].
Additionally, the effectiveness of decision aids may depend on the institutional ability to support and retain human advisors, posing a scaling challenge. Prior work has shown that increased capacity in human outreach programs leads to larger effect sizes and higher power in RCTs [16]. This effect may have contributed to the success of the program at GSU, which has invested significantly in expanding its advising staff and, by the time of the experiment, the student-to-advisor ratio had dropped from 700:1 in 2011 [75] to 239:1, with a 150:1 ratio in the treatment arm as the MAAPS protocol requires [77]. The lowered ratio may be critical to giving advisors the necessary time to consider non-algorithmic context. By contrast, many institutions in the MAAPS experiment did not maintain a 150:1 ratio in the treatment arm due to retention problems, another factor study authors cite for the lack of positive impact at these schools [77]. It is possible that schools adopting algorithmic systems as cost-saving measures to avoid hiring more advisors—placing existing advisors in resource-constrained situations where they feel pressured to follow algorithmic recommendations [47] and cannot develop the skills needed to provide discretionary support—would fail to replicate positive outcomes.
7.4 The Practical and Normative Role of Discretion